# Efficient and Responsible Adaptation of Large Language Models for Robust Top-k Recommendations

### v2

@@ Abstract.
Conventional recommendation systems (RSs) are typically optimized to enhance performance metrics uniformly across all training samples, inadvertently overlooking the needs of diverse user populations. The performance disparity among various populations can harm the model’s robustness to sub-populations due to the varying user properties. While large language models (LLMs) show promise in enhancing RS performance, their practical applicability is hindered by high costs, inference latency, and degraded performance on long user queries. To address these challenges, we propose a hybrid task allocation framework designed to promote social good by equitably serving all user groups. By adopting a two-phase approach, we promote a strategic assignment of tasks for efficient and responsible adaptation of LLMs. Our strategy works by first identifying the weak and inactive users that receive a suboptimal ranking performance by RSs. Next, we use an in-context learning approach for such users, wherein each user interaction history is contextualized as a distinct ranking task. We evaluate our hybrid framework by incorporating eight different recommendation algorithms and three different LLMs – both open and close-sourced. Our results on three real-world datasets show a significant reduction in weak users and improved robustness to subpopulations without disproportionately escalating costs.
@@ 1. Introduction
Recommendation systems (RSs) have become an integral part of numerous online platforms, assisting users in navigating vast amounts of content to relieve information overload (Jacoby, 1984). While Collaborative Filtering based RSs (Papadakis et al., 2022) primarily rely on user-item interactions to predict users’ preferences for certain candidate items, the utilization of language in recommendations has been prevalent for decades in hybrid and content-based recommenders, mainly through item descriptions and text-based reviews (Lops et al., 2011). Furthermore, conversational recommenders (Sun and Zhang, 2018) have highlighted language as a primary mechanism for allowing users to naturally and intuitively express their preferences (Sanner et al., 2023). Deep recommendation models, trained under the Empirical Risk Minimization (ERM) framework, uniformly minimize the loss function across all training samples. However, these models fail to cater to a diverse set of sub-populations, affecting robustness and fairness across user groups. (Geifman and El-Yaniv, 2017; Wang et al., 2023c; Chaney et al., 2018; Beutel et al., 2017; Yao et al., 2021; Zhang et al., 2021; Beutel et al., 2019)Li et al. (2021a) show that active users who have rated many items receive better recommendations on average than inactive users. This inadvertent disparity in recommendations requires careful scrutiny to ensure equitable recommendation experiences for all users (Gunawardana et al., 2012), thereby promoting social good.
@@
Our approach not only aims to improve the technical robustness of RSs but also to enhance social equity by ensuring that under-served populations with limited interaction histories also receive high-quality recommendations. This can lead to better access to educational resources, job opportunities, and health information, thereby contributing to societal well-being. For instance, recommendation systems can improve access to educational resources by personalizing learning experiences, suggesting tailored videos, articles, or exercises (Mahendra, 2024). In the job market, systems like LinkedIn’s engine match job seekers with relevant opportunities, enhancing the job search experience and reducing unemployment (Geyik et al., 2018). Health recommendation systems provide personalized advice on diet, exercise, and treatments, encouraging healthier behaviors and timely medical information (Center for American Progress, 2022). We also incorporate ethical guidelines to ensure responsible adaptation of LLMs, improving robustness and reducing adaptation costs.
@@ 3.1. Problem Formulation
Here, the triplet $d_{mn}=(u_{m},i_{n},r_{mn})$ represents one data point where a user $u_{m}$ provided a rating of $r_{mn}$ to an item $i_{n}$. Now, if a user $u_{m}$ has rated a set of items, then let $[r_{mn}]_{n=1}^{N}$ denote the rating vector consisting of explicit rating values ranging from $1$ to $5$ if a user provided a rating and 00 indicating no rating.
@@ 3.2. Identifying Weak Users
Here, $|R|(n-|R|)$ denotes all possible pairs of relevant and irrelevant items.
@@
Definition 0.
Given dataset $\mathcal{D}$ and a recommendation model $\theta^{r}$, we say that a user $u_{m}$ is extremely weak if the likelihood of $\theta^{r}$ being able to rank the relevant items above the irrelevant items is below $t_{p}$ and the rating vector $[r_{mn}]$ has extremely high sparsity. i.e., above $t_{s}$
[…]
@@ 3.3. Designing Natural Language Instructions for Ranking
Closest to our work, Hou et al. (2024) formalized the recommendation problem as a conditional ranking task considering sequential interaction histories as conditions and uses the items retrieved by traditional RS as $candidate$ items. While we aim to design the conditional ranking tasks, our approach differs significantly from theirs as instead of using LLMs as a re-ranking agent for all users; we instruct LLM with the preferences of weak users preferences (sorted in descending order of decreased preference). This technique is detailed below.
@@
3:  $ranked\_pred\leftarrow\theta^{r}(\mathcal{D}_{train}$)
@@
7:     if $\mathcal{P}(u_{m})<t_{p}$&\&\&& & $\mathcal{S}_{I}(u_{m})$  then
@@ 4.1. Experimental Setup
Hyperparameter Search. We conducted an extensive hyperparameter search for each model to ensure optimal performance. The hyperparameter configurations were chosen based on the model’s characteristics and the specific requirements of our recommendation task. For the BPR (Bayesian Personalized Ranking) model, we focused on tuning the learning rate, as it directly influences the convergence speed and stability of the model. We explored a wide range of learning rates: $[5e\text{-}5,1e\text{-}4,5e\text{-}4,7e\text{-}4,1e\text{-}3,5e\text{-}3,7e\text{-}3]$ to balance the trade-off between fast convergence and maintaining model accuracy. In the case of ItemKNN (Item-based K-Nearest Neighbors), we varied two important hyperparameters: $k$, the number of neighbors considered for each item, and the shrinkage parameter, which controls the weight of the neighbors. The values for $k$ ranged from 10 to 400 ($[10,50,100,200,250,300,400]$), and the shrinkage parameter took values $[0.0,0.1,0.5,1,2]$, as adjusting these parameters helps in fine-tuning the balance between overfitting and generalization. For NeuMF (Neural Matrix Factorization), the hyperparameters include the learning rate, which influences the speed of model optimization, and the architecture of the neural network, specifically the hidden layer sizes. We tested learning rates from $[5e\text{-}7,1e\text{-}6,5e\text{-}6,1e\text{-}5,1e\text{-}4,1e\text{-}3]$ and used a multi-layer perceptron (MLP) structure with hidden sizes of $[64,32,16]$. Dropout probabilities ($[0.0,0.1,0.3]$) were also explored to prevent overfitting by randomly deactivating neurons during training. For NNCF (Neural Collaborative Filtering), we explored a wider range of hyperparameters, as this model is more complex. We varied the number of neighbors ($[20,50,100]$) and the size of the neighbor embeddings ($[64,32]$), which affect how much information from neighboring items is used. Additionally, the number of convolution kernels ($[128,64]$) was adjusted to optimize the feature extraction process, while the learning rate was tuned with values $[5e\text{-}5,1e\text{-}4,5e\text{-}4]$. We also experimented with two methods for aggregating neighborhood information: random and knn, which can influence the model’s ability to capture complex patterns in the data. For BERT4Rec, we focused on the learning rate ($[0.0003,0.0005,0.001,0.003]$), which is crucial for fine-tuning the pre-trained model. We also experimented with different fine-tuning ratios ($[0,0.1,0.5]$) to control the degree of adaptation to the specific task at hand, which can help in leveraging domain-specific knowledge while retaining the pre-trained features. For GRU4Rec, we varied the learning rate ($[0.005,0.001,0.0005,0.0001]$) to determine the optimal step size for model training, and the number of layers ($[1,2]$) to assess the impact of model depth on recommendation quality. Finally, for SASRec, we focused on the learning rate alone, testing values in $[0.0003,0.0005,0.001,0.003,0.005]$, as it plays a critical role in controlling the model’s learning speed and preventing overfitting. This comprehensive hyperparameter search allowed us to explore a wide range of configurations for each model, ensuring that we could optimize each one for the specific characteristics of our datasets and tasks.
@@ 4.2.1. Phase One: Identifying Weak and Strong Users
This analysis also highlights the presence and impact of outliers in the datasets. Outliers, which are users with exceptionally high or low interaction counts, can significantly influence model performance. In sparse datasets, outliers with high interaction counts can skew the results, making it appear that the model performs better than it does for the majority of users. Conversely, outliers with very low interaction counts can highlight the limitations of simpler models like ItemKNN and BPR, which struggle to provide accurate recommendations with minimal data. Our plots show that advanced models like BERT4Rec and GRU4Rec are less affected by outliers due to their ability to generalize from sequential and contextual patterns. This robustness is particularly evident in the Amazon Software and Amazon Video Games datasets, where these models maintain high performance despite the presence of outliers. This instance-by-instance analysis highlights that though on average, sparse users receive poor recommendation results (AUC), not all sparse users experience uniformly poor results. Some sparse users may still receive relatively accurate recommendations depending on specific factors such as their interaction history or the diversity of their preferences. We thus use our definition 5 to identify such users and mark them as weak.
@@ 4.2.2. Second Phase: Evaluation of Our Framework
. We design instructions for the weak users identified in the previous phase using the approach discussed in Section 3. We pass these instructions to three different LLMs (GPT 4, Claude 3.5-haiku, LLama3-70b-instruct) and report AUC values in Table 2. AUC values generally reflect the model’s ability to rank items accurately. A higher AUC indicates better ranking, meaning the model’s ability to recommend relevant items is superior. A model with an AUC near $1.0$ (like SASRec+GPT-4 in Amazon Software) indicates excellent performance in predicting items that a user will engage with. However, values close to $0.5$ imply random recommendations, which is observed in models like ItemKNN for weak users. Our results show that LLMs perform significantly well on these users. It is evident that adding LLMs as re-rankers to various RS algorithms for weak users generally enhances performance. The evaluation is done in terms of AUC across three categories: overall AUC, AUC for weak users, and AUC for non-weak users. From the results, we can see hybrid frameworks, particularly with models like GPT-4, consistently improve the performance for most algorithms, especially in the ”Weak Users” category, where there are notable improvements across ItemKNN, NeuMF, DMF, BPR, and SASRec. The addition of LLMs generally shifts the performance curve upward for these models, as shown in ItemKNN+GPT 4, which shows an improvement in AUC from $0.75100$ (baseline ItemKNN) to 0.80181, reflecting better overall recommendation quality as well as significant improvement for weak users. We emphasize that while the overall AUC improves, we do not pass instructions for non-weak users to LLMs. Thus the overall score remains the same irrespective of the hybrid or non-hybrid framework. Existing post-processing techniques like Li et al. (2021a) show reduced performance on advantaged users to mitigate biases. However, our framework does not affect recommendation quality of non-weak users. We further observe a specific exception in the models involving Claude 3.5-haiku, where adding the model doesn’t improve performance as significantly as GPT-4 or LLama3-70b-instruct for many algorithms, such as ItemKNN+Claude 3.5-haiku, where there is a drop in performance for weak users (AUC $0.57915$) compared to ItemKNN (AUC $0.33172$) and ItemKNN+GPT 4 (AUC $0.59954$). Additionally, the BERT4Rec and GRU4Rec models show robust performance improvement across both weak and non-weak users when paired with GPT-4, highlighting that the synergy between these models is significant for both categories, with improvements in AUC reaching as high as $0.88553$ and $0.88294$, respectively. Across all models, weak users consistently show lower AUC scores compared to non-weak users. This is expected because weak users typically have fewer interactions, leading to poorer model training and prediction accuracy. However, LLM-augmented models tend to reduce this performance gap, as seen in the GPT-4-augmented models, which show substantial improvement in handling weak users. For instance, SASRec+GPT-4 achieves AUC of $0.6699$ for weak users in ML1M, which is significantly better than its non-LLM version. Furthermore, while Claude 3.5-haiku shows some promise, its impact is not as profound, which could indicate that further tuning and integration of the specific LLM may be needed to achieve comparable results to GPT-4 and LLama3-70b-instruct. Overall, the integration of LLMs improves the handling of weak users, reducing the gap between weak and non-weak user performance.
We further conducted significance testing using paired t-test and observed highly significant test statistics across all scenarios, with t-statistics ranging from $-24.60$ to $-60.58$. Correspondingly, the p-values are exceedingly small, often approaching zero i.e. less than $0.05$(e.g., $1.905\times 10^{-304}$ and $0.0$), far below any conventional significance threshold (e.g., $0.05$ or $0.01$). These results strongly suggest rejecting the null hypothesis in each case, indicating substantial evidence against it. The extreme values of the t-statistics, combined with near-zero p-values, reflect robust statistical significance, likely attributable to large effect sizes.
Table 2. Overall comparison in terms of ranking quality measured using AUC across various recommendation models (ItemKNN, NeuMF, NNCF, DMF, BPR, BERT4Rec, GRU4Rec, and SASRec) in comparison to their usage within our framework along with one open-sourced LLM (Claude-3.5-haiku) and two close-sourced LLMs (GPT-4 and LLama-3-70B-insturct).

### v3

@@ Abstract.
Conventional recommendation systems (RSs) often optimize for aggregate accuracy, inadvertently underserving users with sparse interaction histories. While large language models (LLMs) offer strong zero-shot capabilities for ranking, their widespread adoption in RS pipelines is hindered by high computational cost, inference latency, and limited scalability. As a result, most existing works evaluate LLMs on a small, randomly sampled subset of users, raising questions about generalizability and real-world applicability in recommendation scenarios. To address these challenges, we propose a hybrid task allocation framework that responsibly allocates ranking tasks between traditional RSs and LLMs to enhance robustness and efficiency. Our strategy works by first identifying weak and inactive users that receive suboptimal ranking performance from RSs. Next, we use an in-context learning approach for such users, wherein each user’s interaction history is contextualized as a distinct ranking task. We evaluate our hybrid framework by incorporating eight different recommendation algorithms and three different LLMs – both open- and closed-sourced. Our results show that the proposed framework significantly reduces the number of weak users (~12%) while maintaining cost-effectiveness through targeted LLM usage.
@@ 1. Introduction
Recommendation systems (RSs) have become an integral part of numerous online platforms, assisting users in navigating vast amounts of content to relieve information overload (Jacoby, 1984). While Collaborative Filtering based RSs (Papadakis et al., 2022) primarily rely on user-item interactions to predict users’ preferences for certain candidate items, the utilization of language in recommendations has been prevalent for decades in hybrid and content-based recommenders, mainly through item descriptions and text-based reviews (Lops et al., 2011). Furthermore, conversational recommenders (Sun and Zhang, 2018) have highlighted language as a primary mechanism for allowing users to naturally and intuitively express their preferences (Sanner et al., 2023). Deep recommendation models, trained under the Empirical Risk Minimization (ERM) framework, uniformly minimize the loss function across all training samples. However, these models fail to cater to a diverse set of sub-populations, affecting robustness and fairness across user groups (Geifman and El-Yaniv, 2017; Wang et al., 2023c; Chaney et al., 2018; Beutel et al., 2017; Yao et al., 2021; Zhang et al., 2021; Beutel et al., 2019)Li et al. (2021a) show that active users who have rated many items receive better recommendations on average than inactive users. This performance gap motivates the need for robust RS frameworks that can generalize more robustly across different levels of user activity (Gunawardana et al., 2012).
@@
Our approach aims to improve the robustness and adaptability of LLM-based recommendation systems by targeting users with sparse interaction histories, for whom traditional RSs often perform poorly. These users are prevalent across real-world applications where engagement may be infrequent, episodic, or context-dependent. For instance, educational platforms rely on recommendation engines to personalize learning resources such as videos and practice exercises (Mahendra, 2024), but many learners engage sporadically, making accurate personalization challenging. Similarly, job recommendation systems like LinkedIn’s talent engine (Geyik et al., 2018) must surface relevant opportunities for users with limited prior activity, especially those entering the platform for the first time. In healthcare applications, users may interact infrequently with health tracking systems, yet require timely and accurate suggestions related to lifestyle or treatment options (Center for American Progress, 2022). By selectively applying LLMs to these cases, our framework delivers higher-quality recommendations where RSs struggle, while also minimizing unnecessary model overhead. This leads to a more efficient and scalable hybrid system for diverse recommendation tasks. Unlike prior works that uniformly evaluate LLMs on randomly sampled user sets, our contribution lies in a strategic task allocation framework that identifies exactly which users are likely to benefit from LLM augmentation. This improves both computational efficiency and ranking robustness, while avoiding unnecessary inference costs. We focus on how existing RS and LLM systems can be orchestrated more effectively under real-world constraints.
@@ 3.1. Problem Formulation
Here, the triplet $d_{mn}=(u_{m},i_{n},r_{mn})$ represents one data point where a user $u_{m}$ provided a rating of $r_{mn}$ to an item $i_{n}$. Now, if a user $u_{m}$ has rated a set of items, then let $r_{mn}]_{n=1}^{N}$ denote the rating vector consisting of explicit rating values ranging from $1$ to $5$ if a user provided a rating and 0 indicating no rating.
@@ 3.2. Identifying Weak Users
Here, |R|(n|R|)|R|(n-|R|) denotes all possible pairs of relevant and irrelevant items.
@@
Definition 1. Given dataset $\mathcal{D}$ and a recommendation model $\theta^{r}$, we say that a user $u_{m}$ is extremely weak if the likelihood of $\theta^{r}$ being able to rank the relevant items above the irrelevant items is below $t_{p}$, and the user’s interaction vector $\mathbf{r}_{m*}$ exhibits extremely high sparsity i.e., exceeds $t_{s}$.
(5) $\mathcal{P}(u_{m})\leq t_{p}\quad\text{and}\quad S_{I}(u_{m})>t_{s}$
@@ 3.3. Designing Natural Language Instructions for Ranking
Closest to our work, Hou et al. (2024) formalized the recommendation problem as a conditional ranking task considering sequential interaction histories as conditions and uses the items retrieved by traditional RS as candidatecandidate items. While we aim to design the conditional ranking tasks, our approach differs significantly from theirs as instead of using LLMs as a re-ranking agent for all users; we instruct LLM with the preferences of weak users preferences (sorted in descending order of decreased preference). This technique is detailed below.
@@
[…]
@@
7:  if $\mathcal{P}(u_{m})<t_{p}$ &&\&\& $\mathcal{S}_{I}(u_{m})$ then
@@ 4.1. Experimental Setup
  • Fairness-Aware Baseline: Since the adaptation of LLMs for ranking tasks can be interpreted as a post-hoc re-ranking strategy, we include a well-established fairness-oriented post-hoc baseline for comparison:
    • FairRec (Li et al., 2021a): FairRec is a user-oriented fairness framework that aims to balance recommendation quality across user groups, particularly between advantaged (active) and disadvantaged (inactive) users. It formulates the fairness-constrained re-ranking problem as a 0-1 integer programming task, ensuring that the final ranking optimizes utility while satisfying fairness constraints. FairRec has been shown to significantly improve fairness as performance disparity in active and inactive users.
@@
Hyperparameter Search. We conducted an extensive hyperparameter search for each model to ensure optimal performance. The hyperparameter configurations were chosen based on the model’s characteristics and the specific requirements of our recommendation task. For the BPR (Bayesian Personalized Ranking) model, we focused on tuning the learning rate, as it directly influences the convergence speed and stability of the model. We explored a wide range of learning rates: $5e\text{-}5,1e\text{-}4,5e\text{-}4,7e\text{-}4,1e\text{-}3,5e\text{-}3,7e\text{-}3]$ to balance the trade-off between fast convergence and maintaining model accuracy. In the case of ItemKNN (Item-based K-Nearest Neighbors), we varied two important hyperparameters: $k$, the number of neighbors considered for each item, and the shrinkage parameter, which controls the weight of the neighbors. The values for $k$ ranged from 10 to 400 ($[10,50,100,200,250,300,400]$), and the shrinkage parameter took values $[0.0,0.1,0.5,1,2]$, as adjusting these parameters helps in fine-tuning the balance between overfitting and generalization. For NeuMF (Neural Matrix Factorization), the hyperparameters include the learning rate, which influences the speed of model optimization, and the architecture of the neural network, specifically the hidden layer sizes. We tested learning rates from $5e\text{-}7,1e\text{-}6,5e\text{-}6,1e\text{-}5,1e\text{-}4,1e\text{-}3]$ and used a multi-layer perceptron (MLP) structure with hidden sizes of $[64,32,16]$. Dropout probabilities ($[0.0,0.1,0.3]$) were also explored to prevent overfitting by randomly deactivating neurons during training. For NNCF (Neural Collaborative Filtering), we explored a wider range of hyperparameters, as this model is more complex. We varied the number of neighbors ($[20,50,100]$) and the size of the neighbor embeddings ($[64,32]$), which affect how much information from neighboring items is used. Additionally, the number of convolution kernels ($[128,64]$) was adjusted to optimize the feature extraction process, while the learning rate was tuned with values $5e\text{-}5,1e\text{-}4,5e\text{-}4]$. We also experimented with two methods for aggregating neighborhood information: random and knn, which can influence the model’s ability to capture complex patterns in the data. For BERT4Rec, we focused on the learning rate ($[0.0003,0.0005,0.001,0.003]$), which is crucial for fine-tuning the pre-trained model. We also experimented with different fine-tuning ratios ($[0,0.1,0.5]$) to control the degree of adaptation to the specific task at hand, which can help in leveraging domain-specific knowledge while retaining the pre-trained features. For GRU4Rec, we varied the learning rate ($[0.005,0.001,0.0005,0.0001]$) to determine the optimal step size for model training, and the number of layers ($[1,2]$) to assess the impact of model depth on recommendation quality. Finally, for SASRec, we focused on the learning rate alone, testing values in $[0.0003,0.0005,0.001,0.003,0.005]$, as it plays a critical role in controlling the model’s learning speed and preventing overfitting. This comprehensive hyperparameter search allowed us to explore a wide range of configurations for each model, ensuring that we could optimize each one for the specific characteristics of our datasets and tasks.
@@ 4.2.1. Phase One: Identifying Weak and Strong Users
This analysis also highlights the presence and impact of outliers in the datasets. Outliers, which are users with exceptionally high or low interaction counts, can significantly influence model performance. In sparse datasets, outliers with high interaction counts can skew the results, making it appear that the model performs better than it does for the majority of users. Conversely, outliers with very low interaction counts can highlight the limitations of simpler models like ItemKNN and BPR, which struggle to provide accurate recommendations with minimal data. Our plots show that advanced models like BERT4Rec and GRU4Rec are less affected by outliers due to their ability to generalize from sequential and contextual patterns. This robustness is particularly evident in the Amazon Software and Amazon Video Games datasets, where these models maintain high performance despite the presence of outliers. This instance-by-instance analysis highlights that though on average, sparse users receive poor recommendation results (AUC), not all sparse users experience uniformly poor results. Some sparse users may still receive relatively accurate recommendations depending on specific factors such as their interaction history or the diversity of their preferences. We thus use our definition to identify such users and mark them as weak.
@@ 4.2.2. Second Phase: Evaluation of Our Framework
. We design instructions for the weak users identified in the previous phase using the approach discussed in Section 3.Table 3 presents a comprehensive comparison of AUC scores across multiple recommendation models and datasets, revealing several important trends regarding the effectiveness of LLM-based augmentation and fairness-aware baseline (Li et al., 2021a). Across all model families — including shallow (e.g., ItemKNN, BPR) and deep learning-based approaches (e.g., GRU4Rec, SASRec) — we observe that integrating LLMs, especially GPT-4, leads to consistent and substantial gains in recommendation quality for weak users. For instance, in the ML1M dataset, SASRec+GPT-4 improves AUC for weak users from 0.5109 (base) to 0.6690 — a relative gain of over 15.8%. Similarly, GRU4Rec+GPT-4 shows a dramatic improvement from 0.4021 to 0.6900 on weak users. These gains are not limited to ML1M but are consistently observed in Amazon Video Games and Amazon Software domains, with SASRec+GPT-4 achieving 0.6645 and 0.8191 respectively for weak users. Importantly, performance for non-weak users remains largely unchanged as recommendations for these users are presented via traditonal models.
We also compare our method with fairness-aware baselines such as FairRec-enhanced variants. While models like SASRec+FairRec (0.6485) and BPR+FairRec (0.5963) improve over their base counterparts, they consistently underperform compared to their GPT-4-augmented versions (e.g., SASRec+GPT-4: 0.6690; BPR+GPT-4: 0.6110 on ML1M weak users). Furthermroe, FairRec enhanced versions further reduce the perfomance over non-weak users in order to reduce the performance gap between weak and non-weak users. This trend highlights that our approach offers stronger instance-level adaptation through contextualization rather than relying on group-based fairness post-processing. The consistency of LLM gains across all models and datasets further supports the model-agnostic and domain-general nature of our responsible adaptation framework. Additionally, these results underscore the cost-efficiency of our design: by selectively allocating LLMs to weak users, we avoid blanket inference while still achieving superior fairness outcomes. Thus, our findings affirm the scalability and robustness of our hybrid LLM-RS framework, demonstrating its potential to serve as a practical and principled solution for improving user-side fairness in real-world recommendation systems.
To assess the robustness of our improvements, we perform paired two-tailed t-tests comparing our hybrid approach with each baseline across all user-level metric (AUC). For each user, we compute the metric difference between the hybrid model and a baseline, resulting in a paired sample of differences. We then compute the t-statistic and corresponding p-value for each model-dataset combination. The t-statistics range from 24.60-24.60 to 60.58-60.58, with p-values consistently below $0.05$, often approaching zero (e.g., $1.905\times 10^{-304}$ and $0.0$). These results indicate statistically significant improvements, providing strong evidence against the null hypothesis. The consistently extreme t-statistics and vanishingly small p-values reflect large effect sizes.
4.2.3. Cost Analysis
We compare our hybrid framework against a naive baseline in which all users are routed through the LLM for ranking. Table 2 reports the total token usage and estimated latency savings across datasets. Full LLM Tokens refers to the total number of input/output tokens consumed when all user ranking tasks are handled by the LLM, including both prompt and response tokens. Hybrid Tokens represents the total token usage when only weak users (identified using AUC and sparsity thresholds) are routed to the LLM, while strong users are handled by the base RS model—incurring no LLM token cost. Latency Savings is computed as the percentage reduction in average end-to-end response time when using the hybrid approach instead of a full LLM pipeline. Latency is measured from the time a prompt is sent to the LLM API to when the response (ranked list) is received. We estimate average per-user latency based on practical assumptions: LLM inference takes approximately 8 seconds per user, while base RS models operate at roughly 0.02 seconds per user. Let $N$ be the number of users and $W$ the number of weak users (estimated at 30% of $N$). Then, total latency for the full LLM pipeline is $N\times 8$, and for the hybrid setup it is $W\times 8+(N-W)\times 0.02$. The latency savings is computed as:
$\left(1-\frac{\text{Hybrid Latency}}{\text{Full LLM Latency}}\right)\times 100.$
To evaluate token usage, we estimate counts using OpenAI’s GPT-4 tokenizer, assuming prompts include 20 training items and 10 test items per user. Latency estimates are based on inference time averaged over 500 randomly sampled users per dataset. Our hybrid approach reduces total token usage by over 60% and reduces latency by around 70% across all datasets, without compromising overall AUC performance, making it well-suited for scalable deployment.
Table 2. Comparison of total token usage and estimated latency savings between full-LLM and hybrid approach.
Dataset Full LLM Tokens Hybrid Tokens Latency Savings
ML-1M 2.4M 790K 70%
Amazon Software 2.1M 700K 69.83%
Amazon Video Games 1.1M 307K 69.86%
Table 3. Overall comparison in terms of ranking quality measured using AUC across various recommendation models (ItemKNN, NeuMF, NNCF, DMF, BPR, BERT4Rec, GRU4Rec, and SASRec), their LLM-augmented versions (with GPT-4, LLaMA, and Claude), and their fairness-enhanced extensions using FairRec. Results are reported for all users, as well as broken down into weak and non-weak user groups, across three datasets (ML1M, Amazon Video Games, and Amazon Software).